activity preceding optimistic probability pessimistic
a -- 7 9 14
b a 2 2 8
c a 8 12 16
d a 3 5 10
e b 4 6 8
f b 6 8 10
g c,f 2 3 4
h d 2 2 8
i h 6 8 16
j g,i 4 6 14
k e,j 2 2 5

a. What is the estimated time of completing the project if the critical path is A-D-H-I-J-K?

b. calculate the project variance?

c. Compute the project standard deviation?

d. Find the probability that the project will be completed in less than the project completion time?
Answer #1

Answer:

Activity

Optimistic time-a

Expected completion time-m

Pessimistic time-b

Expected time= (a+4*m+ b)/6

Variance, (sigma)^2= (b-a/6)^2

A

7

9

14

9.50

1.36

B

2

2

8

3.00

1.00

C

8

12

16

12.00

1.78

D

3

5

10

5.50

1.36

E

4

6

8

6.00

0.44

F

6

8

10

8.00

0.44

G

2

3

4

3.00

0.11

H

2

2

8

3.00

1.00

I

6

8

16

9.00

2.78

J

4

6

14

7.00

2.78

K

2

2

5

2.50

0.25

Answer a:

the project is expected to take 36.50 days along critical path ADHIJK

Answer b:

total project variance= 9.53 (sum of the variance of activities on critical path)

Answer c:

standard deviation= sqrt(variance)= sqrt(9.53)= 3.08

step 1

we will find the variance of the tasks which lie on critical path ADHIJK

step 2

mean project time (u) of the critical path is=

36.50

step 3

Required completion time is 36.5 weeks

step 4

standard deviation= sqrt(variance)= sqrt(9.53)=

3.087

step 5

because Z= (given completion time- u)/standard deviation=

0

step 6

P(z<0)=

0.5000

50.00%

Answer d:

probability= 50%